# Koala-36M : A Large-scale Video Dataset Improving Consistency between Fine-grained Conditions and Video Content

### v1

@@ 4.3.1 New criteria and human evaluation
We have defined a new annotation criterion that assigns a score reflecting whether a video is suitable as training data for video generation models. This criterion primarily considers the following aspects of video quality: Dynamic Quality: A high-quality video should exhibit good dynamics, which are evaluated based on two factors: the extent of subject movement and the temporal stability of the motion. The motion area in the video should cover more than 30% of the frame; otherwise, the score of the video will be decreased for insufficient dynamics. Temporal stability considers the camera movement; non-professional videographers often produce videos with irregular and significant shaking. We decrease the scores of such videos to distinguish them from professional works. Static Quality: Each frame of a high-quality video should have rich subject details, reasonable composition, aesthetic appeal, clear and distinct subjects, and saturated colors. Although this metric may involve some subjectivity, it is crucial for assessing the overall visual quality. Video Naturalness: We prefer videos that are natural and unprocessed. Special effects, transitions, subtitles, and logos can introduce biases in the original distribution of videos, making it harder for generation models to learn. Additionally, we consider the safety of the video content, rejecting videos with political, terrorist, violent, pornographic, gory, or otherwise disturbing content. In order to reduce the bias between the labeled scores and the true scores, each video is labeled by 8 experts and subjected to a bias elimination process, as described in the App A.2.
@@
Figure 7: The pipeline of Training Suitability Assessment Network.
4.3.2 Training Suitability Assessment Network
As shown in Fig. 7, we propose a Training Suitability Assessment Network, which takes videos and sub-metrics as input, and outputs a single value called Video Training Suitability Score (VTSS). Corresponding to the aforementioned annotation criteria, our network is divided into dynamic and static branches. Additionally, we retain various data labels from traditional data filtering strategies and pass this extra information to the network model as a new branch. For the features of different branches, the 3D Swin Transformer is used as the backbone for the dynamic branch, while the ConvNext network serves as the backbone for the static branch. To integrate the features from different branches, we propose a Weight Cross-Gating Block (WCGB) to incorporate the information from the label branch into the other two branches. Since the label branch inherently reflects various characteristics of the video, which are related to both dynamic and static features, we use label features to enhance the dynamic and static features. Given that different video labels focus on dynamic and static aspects to varying degrees, we learn a fusion weight to adjust the proportion of label features integrated with the two types of video features.
After training Training Suitability Assessment Network on the human-aligned dataset, we employ it to predict Video Training Suitability Score (VTSS) for all videos, and obtain the score distribution as shown in Fig. 6. Since the VTSS distribution can roughly be divided into two Gaussian distributions, we simply chose the decomposition value 2.5 as the VTSS threshold. Based on this threshold, we filtered out a dataset containing a total of 36 million video clips with corresponding captions. And we name the dataset as Koala-36M , which is the final dataset we present.
@@ 4.4 Metrics conditioning
In previous pipelines, data metrics are simply used for data filtering. Meanwhile, the quality of the filtered data still varied, making it difficult for the model to distinguish between high-quality and low-quality data. To address this issue, we propose a more fine-grained conditioning method to incorporate quality information of different videos into the generation model during training, leading to better consistency between conditions and video content. During inference, this method also enables fine-grained control over the generated videos.
@@
Specifically, during video diffusion training, we first encode data metrics such as motion score, aesthetic score, and clarity score into frequency embeddings. Subsequently, frequency embeddings are passed through an MLP to obtain multiple embeddings, which are then directly added to the timestep embeddings and incorporated into the transformer block using Adaptive Layer Normalization (AdaLN). This method has two main advantages. First, it does not increase the computational load of the diffusion model. Second, compared to adding conditions in captions like Open-sora (Zangwei et al., 2024), it allows for more precise control by being more sensitive to numerical scores, and posses a stronger ability to decouple control over different metrics. During the inference stage, we can set different feature scores, such as setting all scores to the highest value, to generate high-quality videos.
5 EXPERIMENTS
5.1 Experiment Setting
To validate he superiority of Koala-36M dataset and the effectiveness of our data processing pipeline, we train the same generation model from scratch on different datasets for comparison. Our text-to-video base model is based on a 3D attention-like Sora structure (Brooks et al., 2024), and the VAE employs a causal convolution-based 3D VAE. Since the training was done from scratch, we set the video duration to 2 seconds and the resolution to 256x256 for faster convergence. All models are trained on their respective datasets passing through 140M data samples in total. To evaluate the performance of generation models, we conduct a comprehensive evaluation on the public benchmark VBench (Huang et al., 2023). Due to the domain gap between the captions provided by VBench and training set, we performed prompt expansion on the captions in VBench.
5.2 Quantitative Results
@@
Koala-all 0.4808 0.2105 0.9335 0.9668 0.9857 0.9855 0.4222 0.5535 0.5453 0.1154
Koala-37M (manual threshold) 0.4683 0.2135 0.9388 0.9664 0.9810 0.9870 0.4028 0.5422 0.4858 0.1099
@@
Koala-all 0.5180 0.8958 0.2168 0.1630 0.1971 0.1881 0.7758 0.4668 0.7140
Koala-37M (manual threshold) 0.4700 0.9128 0.1978 0.1589 0.2003 0.1893 0.7704 0.4548 0.7073
@@ 5.3 Qualitative Results
We visualize the generated videos on VBench’s prompts in Fig. 10. The generation model achieve the optimal performance on Koala-36M , with both the best video quality and text-video consistency. Koala-36M outperform the larger Panda-70M dataset with only 36M data, indicating that our data quality far exceeds that of Panda-70M. See A.5 for more video generation results.
Figure 10: Qualitative results of text-to-video generation. We train the same generation model from scratch on different datasets for comparison. The generation model achieve the optimal performance on Koala-36M , with both the best video quality and text-video consistency.
@@ 5.4 Ablation Experiments
We conduct extensive ablation experiments to demonstrate the superiority of our dataset and the entire pipeline. Specifically, we performed ablation experiments on different data processing and training strategies, divided into the following groups: (1) Panda-70M: baseline. (2) Koala-all: All 48M data after video splitting and captioning. (3) Koala-37M: manually filtered data from Koala-all using multiple thresholds. (4) Koala-36M: filtered dataset from Koala-all using VTSS. (5) Koala-all-condition: Koala-all with metrics conditions. (6) Koala-36M-condition: Koala-36M with metrics conditions.
Data Processing. Comparing the results of training from Panda-70M and Koala-all in Tab. 2 and Fig. 10, we find that Koala-all produce better results, especially in temporal quality, such as subject consistency, background consistency and temporal flickering. This indicates that our newly proposed re-splitting algorithm can more accurately segment transitions, reducing semantic inconsistencies between video segments. Additionally, our recaptioning algorithm provided more detailed video descriptions, making it easier for the model to learn the relationship between visual and textual information. To further demonstrate the superiority of our splitting and captioning methods, we conducted extensive comparative experiments, detailed in the App. A.1.
Data Filtering. Comparing the results of training from Koala-all&Koala-36M and Koala-all-condition&Koala-36M-condition, we find that the results from the latter one perform better than that from the former datasets. This indicates that filtering out low-quality data and retaining high-quality data are necessary to prevent the model from learning biased distributions from low-quality data. In addition, comparing the results of training from Koala-46&Koala-36M, it can be concluded that our filtering method based on single VTSS results in better filtering performance, when more high-quality data and less low-quality data being retained. Extensive ablation experiments of Training Suitability Assessment Network are conducted in the App. A.3.
Metrics conditions. Comparing the results of training from Koala-36M&Koala-36M-condition, the generation model shows significant improvements in video quality, when metrics conditions are injected into it. This indicates that guiding model training using sub-metrics is necessary, as it helps the model implicitly model the importance of different data. In addition, we compare our AdaLN-based injection method with text-encoder based method (Zangwei et al., 2024) in App. A.4 Fig. 13. It can be discovered that our injection method has more precise control and stronger ability to decouple control over different metrics, when the style of videos transfer with the motion score.
@@ 6 Conclusion
Limitations. Despite all the strength above, Koala-36M is still insufficient to support the training of an extremely large video generation model with over 1B parameters. A larger-scale datasets need to be further collected and processed. Meanwhile, the performance, generalization, and scaling laws of generation models on high-quality datasets need further exploration.
Appendix A Appendix
A.1 Effectiveness of video splitting methods
To validate the accuracy and efficiency of our proposed Color-Struct SVM (CSS) for scene transition detection, we conduct the following experiments.
We annotate transitions in 10,000 video clips, creating a test set (approximately half of the videos contain transitions). We then apply our proposed method and open-source methods to detect transitions in the test set, recording the precision and recall of the detections. The open-source method is primarily based on Pyscenedetect (Castellano, ), and we test two versions: one that detects transitions based solely on HSL (Hue, Saturation, Lightness) and another that uses both HSL and edge detection. The experimental results are shown in the Tab. 3. It can be observed that our transition detection algorithm outperforms the two pyscenedetect-based methods in terms of both precision and recall. Notably, our algorithm achieves a high recall rate, indicating that it rarely misses transitions in videos.
@@ A.1 Effectiveness of video splitting methods
On the other hand, we compare the runtime efficiency of our method with that of the open-source algorithms. We record the CPU runtime of our algorithm and other open-source algorithms at different resolutions, with the experimental results shown in Tab 4. We find that at a resolution of 256x256, our method performs comparably to other methods. However, as the video resolution increases, our method becomes significantly faster than the other methods (Fig 11).
@@
Figure 11: Time Consumption for Different Resolutions and Methods. Our method is faster than the others at higher video resolutions.
A.2 Elimination of Deviations between True Scores and Labeled Scores
After establishing the criteria, we randomly sample a batch of data and have it annotated by trained experts, with each video being scored by eight experts on a scale of 1 to 5. To ensure that the annotations closely reflect the true suitability scores, we need to address two types of errors: Individual Preference Bias: As shown in the Fig. 12(a), we visualize the violin plots of scores given by different experts. The expert on the left tends to give lower scores, while the expert on the right tends to give higher scores. These individual preferences can cause the final scores of some videos to be lower or higher than their actual values. Therefore, we standardize the scores of each expert and then scaled them using the mean and variance of the overall scores to eliminate the bias introduced by different experts. From the figure, it can be seen that the scores processed through our normalization and rescaling methods align more closely with the overall score distribution. Label Fluctuation Bias: As shown in the Fig. 12(b), each video is annotated by eight experts, and different experts may assign different scores due to varying interpretations of the criteria. This leads to label fluctuations. We use the mean score to reduce the error caused by these fluctuations.
@@ A.2 Elimination of Deviations between True Scores and Labeled Scores
A.3 Ablation Experiments of Training Suitability Assessment Network
Table 5: Performance Metrics for Different Combinations of Video, Image, and Feature
@@ A.3 Ablation Experiments of Training Suitability Assessment Network
We conduct comprehensive ablation experiments on our Training Suitability Assessment Network. The experimental results are shown in Tab 5 . The baseline model utilizes only dynamic features. Adding the static branch enables the model to capture more static information, thereby improving overall performance. The inclusion of the feature branch allows the model to leverage additional label information, further enhancing its performance. The WCGB module integrates label information with dynamic and static features through a cross-gating mechanism, achieving optimal performance. Each module addition significantly boosts the model’s performance.
Combining dynamic and static branches allows the model to capture both types of information. The feature branch utilizes label information for further improvement. The WCGB module optimizes feature integration, achieving the best results.
A.4 Comparison of results from different metrics conditions

### v2

@@ 4.3.1 New criteria and human evaluation
We have defined a new annotation criterion that assigns a score reflecting whether a video is suitable as training data for video generation models. This criterion chiefly takes into account the following dimensions of video quality: Dynamic Quality: A high-quality video should exhibit good dynamics, which are evaluated based on two factors: the extent of subject movement and the temporal stability of the motion. The motion area in the video should cover more than 30% of the frame; otherwise, the score of the video will be decreased for insufficient dynamics. Temporal stability pertains the camera movement; non-professional videographers often produce videos with irregular and significant shaking. We decrease the scores of such videos to distinguish them from professional works. Static Quality: Each frame of a high-quality video should have rich subject details, reasonable composition, aesthetic appeal, clear and distinct subjects, and vividly saturated colors. Although this metric may involve some subjectivity, it is crucial for assessing the overall visual quality. Video Naturalness: We prefer videos that are natural and unprocessed. Special effects, transitions, subtitles, and logos can introduce biases in the original video distribution , making it harder for generation models to learn. Additionally, we consider the safety of the video content, rejecting videos with political, terrorist, violent, gory, or otherwise disturbing content. In order to reduce the bias between the labeled scores and the true scores, each video is labeled by 8 experts and subjected to a bias elimination process in the App B.
4.3.2 Training Suitability Assessment Network
As shown in Fig. 6, we propose a Training Suitability Assessment Network, which takes videos and sub-metrics as input, and outputs a single value called Video Training Suitability Score (VTSS). Aligned with the previously mentioned annotation criteria, our network is structured into dynamic and static branches. Furthermore, we maintain diverse data labels from conventional data filtering approaches and introduce this supplementary information to the network model as an additional branch. For the features of different branches, 3D Swin Transformer is employed as the backbone for the dynamic branch, while the ConvNext network for the static branch. To integrate the features from different branches, we propose a Weight Cross-Gating Block (WCGB) to incorporate the information from the label branch into the other two branches. Since the label branch inherently reflects various characteristics of the video, which are related to both dynamic and static features, we use label features to enhance the dynamic and static features. Given that different video labels focus on dynamic and static aspects to varying degrees, we learn a fusion weight to adjust the proportion of label features integrated with the two types of video features.
Figure 6: Training Suitability Assessment Network.
After training Training Suitability Assessment Network on the human-aligned dataset, we employ it to predict Video Training Suitability Score (VTSS) for all videos, and obtain the score distribution as shown in Fig. 7. Since the VTSS distribution can roughly be divided into two Gaussian distributions, we simply chose the decomposition value 2.5 as the VTSS threshold. Based on this threshold, we filtered out a dataset containing a total of 36 million video clips with corresponding captions. We designate this dataset as Koala-36M , which is the final dataset we are presenting.
@@ 4.4 Metrics conditioning
In previous pipelines, data metrics are simply used for data filtering. Meanwhile, the quality of the filtered data still varied, making it difficult for the model to distinguish between high-quality and low-quality data. To address this issue, we propose a more fine-grained conditioning method to incorporate quality information of different videos into the generation model during the training phase, leading to better consistency between conditions and video content. During the inference stage, this method also enables fine-grained control over the generated videos.
Specifically, during video diffusion training, we first encode data metrics such as motion score, aesthetic score, and clarity score into frequency embeddings. Subsequently, frequency embeddings are passed through an MLP to obtain multiple embeddings, which are then directly added to the timestep embeddings and incorporated into the transformer block using Adaptive Layer Normalization (AdaLN). This method offers two main advantages. First, it does not increase the computational load of the diffusion model. Second, compared to adding conditions in captions like Open-sora [39], it supports more precise control with higher sensitive to numerical scores, and posses a stronger ability to decouple control over different metrics. During the inference, we can set different feature scores, such as setting all scores to the highest value, to generate high-quality videos.
@@
Koala-w/o TSA 0.4808 0.2105 0.9335 0.9668 0.9857 0.9855 0.4222 0.5535 0.5453 0.1154
Koala-37M-manual 0.4683 0.2135 0.9388 0.9664 0.9810 0.9870 0.4028 0.5422 0.4858 0.1099
@@
Koala-w/o TSA 0.5180 0.8958 0.2168 0.1630 0.1971 0.1881 0.7758 0.4668 0.7140
Koala-37M-manual 0.4700 0.9128 0.1978 0.1589 0.2003 0.1893 0.7704 0.4548 0.7073
@@
5 Experiments
5.1 Experiment Setting
To validate he superiority of Koala-36M dataset and the effectiveness of our data processing pipeline, we train the same generation model from scratch on different datasets for comparison. Our text-to-video base model is based on a Sora-like structure [4] with 3D-full attention transformer block. And each basic transformer block includes 2D self-attention, 3D self-attention, and text cross-attention. We use T5 for text embedding and 3D causal VAE for video compression. Since the training was done from scratch, we set the video duration to 2 seconds and the resolution to 256x256 for faster convergence. We train models on 80G A100 GPUs with a batch size of 32 and a learning rate of 0.0001. All models are trained on their respective datasets passing through 140M data samples in total. To evaluate the performance of generation models, we conduct a comprehensive evaluation on the public benchmark VBench [13]. Due to the domain gap between the captions provided by VBench and training set, we performed prompt expansion on the captions in VBench.
5.2 Quantitative Results
@@ 5.3 Qualitative Results
We visualize the generated videos on VBench’s prompts in Fig. 10. The generation model achieve the optimal performance on Koala-36M , with both the best video quality and text-video consistency. Koala-36M outperform the larger Panda-70M dataset with only 36M data, indicating that our data quality far exceeds that of Panda-70M. See G for more video generation results.
Figure 10: Qualitative results of text-to-video generation. We train the same generation model from scratch on different datasets for comparison. The generation model achieve the optimal performance on Koala-36M , with better video quality and text-video consistency.
@@ 5.4 Ablation Experiments
We conduct extensive ablation experiments to demonstrate the superiority of our dataset and the entire pipeline. Specifically, we performed ablation experiments on different data processing and training strategies, divided into the following groups: (1) Panda-70M: baseline. (2) Koala-w/o TSA: All 48M data without filtering after video splitting and captioning. (3) Koala-37M-manual: data filtered from all 48M data by manually multiple thresholds instead of VTSS. (4) Koala-36M: filtered dataset from Koala-w/o TSA using VTSS. (5) Koala-w/o TSA-condition: All 48M data without filtering but with metrics conditions. (6) Koala-36M-condition: Koala-36M with metrics conditions.
Data Processing. Comparing the results of training from Panda-70M and Koala-w/o TSA in Tab. 2 and Fig. 10, we find that Koala-w/o TSA produce better results, especially in temporal quality, such as subject consistency, background consistency and temporal flickering. This indicates that our newly proposed re-splitting algorithm can more accurately segment transitions, reducing semantic inconsistencies between video segments. Additionally, our recaptioning algorithm provided more detailed video descriptions, making it easier for the model to learn the relationship between visual and textual information. To further demonstrate the superiority of our splitting and captioning methods, we conducted extensive experiments in the App. A.
Data Filtering. Comparing the results of training from Koala-w/o TSA and Koala-36M, Koala-w/o TSA-condition and Koala-36M-condition, we find that the results from the latter one perform better than that from the former datasets. This indicates that filtering out low-quality data and retaining high-quality data are necessary to prevent the model from learning biased distributions from low-quality data. In addition, comparing the results of training from Koala-37M-manual and Koala-36M, it can be concluded that our filtering method based on single VTSS results in better filtering performance, when more high-quality data and less low-quality data being retained. Extensive experiments of Training Suitability Assessment Network are in App. C.
Metrics conditions. Comparing the results of training from Koala-36M and Koala-36M-condition, the generation model shows significant improvements in video quality, when metrics conditions are injected into it. This indicates that guiding model training using sub-metrics is necessary, as it helps the model implicitly model the importance of different data. In addition, we compare our AdaLN-based injection method with text-encoder based method [39] in App. F Fig. 13. It can be discovered that our injection method has more precise control and stronger ability to decouple control over different metrics, when the style of videos transfer with the motion score.
@@ 6 Conclusion
Limitations. Despite all the strength above, Koala-36M is still insufficient to support the training of an extremely large video generation model with over 1B parameters. A larger-scale datasets need to be further collected and processed, which is remained as the future work.
Supplementary Material
Appendix A Effectiveness of video splitting methods
To validate the accuracy and efficiency of our proposed Color-Struct SVM (CSS) for scene transition detection, we conduct the following experiments. We annotate transitions in 10,000 video clips, creating a test set (approximately half of the videos contain transitions). We then apply our proposed method and open-source methods to detect transitions in the test set, recording the precision and recall of the detections. The open-source method is primarily based on Pyscenedetect [6], and we test two versions: one that detects transitions based solely on HSL (Hue, Saturation, Lightness) and another that uses both HSL and edge detection. As shown in the Tab. 3, it can be observed that our transition detection algorithm outperforms the two pyscenedetect-based methods in terms of both precision and recall. Notably, our algorithm achieves a high recall rate, indicating that it rarely misses transitions in videos.
@@ Appendix A Effectiveness of video splitting methods
On the other hand, we compare the runtime efficiency of our method with that of the open-source algorithms. We record the CPU runtime of our algorithm and other open-source algorithms at different resolutions, with the experimental results shown in Tab. 4. We find that our method performs comparably to other methods at 256 resolution. However, as the video resolution increases, our method becomes significantly faster than the other methods (Fig. 11).
Figure 11: Time Consumption for Different Resolutions. Our method is faster than the others at higher video resolutions.
@@
Appendix B Elimination of Deviations between True Scores and Labeled Scores
After establishing the criteria, we randomly sample 200k videos and have it annotated by trained experts, with each video being scored by eight experts on a scale of 1 to 5. To ensure that the annotations closely reflect the true suitability scores, we need to address two types of errors: Individual Preference Bias: As shown in the Fig. 12(a), we visualize the violin plots of scores given by different experts. The expert on the left tends to give lower scores, while the expert on the right tends to give higher scores. These individual preferences can cause the final scores of some videos to be lower or higher than their actual values. Therefore, we standardize the scores of each expert and then scaled them using the mean and variance of the overall scores to eliminate the bias introduced by different experts. From the figure, it can be seen that the scores processed through our normalization and rescaling methods align more closely with the overall score distribution. Label Fluctuation Bias: As shown in the Fig. 12(b), each video is annotated by eight experts, and different experts may assign different scores due to varying interpretations of the criteria. This leads to label fluctuations. We use the mean score to reduce the error caused by these fluctuations.
@@ Appendix B Elimination of Deviations between True Scores and Labeled Scores
Appendix C Ablation Study of Training Suitability Assessment Network
We conduct comprehensive ablation experiments on our Training Suitability Assessment Network. The experimental results are shown in Tab. 5. The baseline model utilizes only dynamic features. Adding the static branch enables the model to capture more static information, thereby improving overall performance. The inclusion of the feature branch allows the model to leverage additional label information, further enhancing its performance. The WCGB module integrates label information with dynamic and static features through a cross-gating mechanism, achieving optimal performance. Each module addition significantly boosts the model’s performance. Combining dynamic and static branches allows the model to capture both types of information. The feature branch utilizes label information for further improvement. The WCGB module optimizes feature integration, achieving the best results. In Tab. 6, we further supplement the results of Dover to demonstrate the superiority of our TSA module.
Table 5: Performance of Different Combinations
@@ Appendix C Ablation Study of Training Suitability Assessment Network
Table 6: Comparison between Dover and TSA.
[…]
FastVQA 0.8684 0.8580 0.7027 0.4644
Dover 0.8554 0.8506 0.6788 0.6497
Ours 0.8974 0.8868 0.7406 0.4099
Appendix D Further Analysis on Data Filtering
We claim existing methods neglect the joint distribution of sub-metrics, resulting in inaccurate thresholds in Sec. 4.3. We demonstrate our claim with two conclusions. (1) There exists joint distribution between sub-metrics. As shown in Tab. 7, we select three sub-metrics and calculate the pairwise correlation coefficients on the unfiltered data, finding that the different sub-metrics are not independent of each other. (2) Due to the slight deviation of the sub-metrics threshold from the optimal values, the filtering errors accumulate on the errors of each sub-metric, resulting in a larger overall filtering error. As shown in Tab. 8, the amount of incorrectly filtered data increases as the number of sub-metrics with inaccurate threshold increases.
Table 7: Correlation coefficients between sub-metrics.
Correlation coefficients (Clarity, Aesthetic) (Clarity, Motion) (Motion, Aesthetic)
Pearson correlation 0.3774 -0.4028 -0.2515
Spearman’s rank correlation 0.3732 -0.4324 -0.2347
Table 8: The incorrectly filtered data with the increasing number of inaccurate sub-metrics thresholds.
Sub-metrics threshold with deviation (+10%) (Clarity) (Clarity, Aesthetic) (Clarity, Motion, Aesthetic)
Incorrectly filtered data / all data 250K/48M 290K/48M 340K/48M
Appendix E More Quantitative Results
As shown in Tab. 9, we further pretrain the same model on the HD-VG-130M dataset and evaluate its performance on VBench and additional metrics, such as FVD score. Our dataset outperforms Panda-70M and HD-VG across all metrics. The non-training metrics of other datasets are also presented in Tab. 1 of the main paper. Meanwhile, we further conduct experiments on higher resolution (512) and longer duration (4s), demonstrating the superiority of our dataset.
Table 9: Quantitative results of text-to-video generation
[…]
Panda 256-2s 0.7343 0.3093 0.6493 570.87
HD-VG-130M 256-2s 0.7696 0.4541 0.7065 590.86
Koala-36M (condition) 256-2s 0.7846 0.5919 0.7460 549.79
Panda 256-4s 0.7395 0.4448 0.6806 451.09
Koala-36M (condition) 256-4s 0.7644 0.4646 0.7045 354.79
Panda 512-2s 0.7439 0.3954 0.6742 579.57
Koala-36M (condition) 512-2s 0.7849 0.6495 0.7578 392.26
Appendix F Comparison of results from different metrics conditions